Grade 12th passMechanics

how far does your car moving at 55mi/h travel forward during the 1s of time that you take to look at an accident on the side of direction?

Muhammad faisal
9 Years agoGrade 12th pass

1 Answer

Vikas TU
9 Years ago
For that ones econd the distance travelled would be obvious,
distance = speed x time
              = 55*1000*1/3600
              = 550/360
              = 55/36 
              = 15.277 meter.